# Sensing-Enhanced Channel Estimation for Near-Field XL-MIMO Systems

### v1

@@ IV-B Low-Complexity Time Inversion Algorithm
where $\mathcal{F}_{a}[\cdot]$ denotes an FFT operation with padding length $2^{\lceil\log_{2}a\rceil}$, and $\mathcal{F}^{-1}_{a}[\cdot]$ denotes an inverse FFT (IFFT) operation with trimming length $a$. By resorting to FFT, the computational complexity can be greatly reduced from $\mathcal{O}(N_{\rm BS}G_{x})$ to $\mathcal{O}(P\log P)$, where $P=2^{\lceil\log_{2}G_{x}\rceil}$ is the nearest power-$2$ number larger than $G_{x}$.
0:  The range of near-field region $x_{L}$, $x_{H}$, $y_{L}$, and $y_{H}$, the numbers of grids $G_{x}$ and $G_{y}$, multipaths $L$,666 To obtain the value of $L$, an empirical threshold can be adopted to identify the peaks in $\bm{\Theta}$ that are significantly higher than other ripples.  and antennas $N_{\rm BS}$. Measured power pattern $\bf p$ and reference wave ${\bf e}_{\rm ref}$.
@@ V-A Eigen-Dictionary Design for Sparse Representation
${\bf R}_{\rm BS}$ $\overset{~{}~{}~{}}{=}\mathbb{E}\left[{\bf H}^{H}{\bf H}\right]$ (36)
$\overset{~{}~{}~{}}{=}{\bf H}_{\rm LoS}^{H}{\bf H}_{\rm LoS}+\mathbb{E}\left[{\bf H}_{\rm NLoS}^{H}{\bf H}_{\rm NLoS}\right]$
$\overset{~{}~{}~{}}{=}{\bf H}_{\rm LoS}^{H}{\bf H}_{\rm LoS}+\bm{\Gamma},$
@@
${\bf R}_{\rm BS}{\bf v}_{n}={\bf D}_{\rm BS}^{-1}\left(\overline{\mathbf{R}}_{\rm BS}+{\bm{\Gamma}}\right){\bf D}_{\rm BS}{\bf v}_{n}=\lambda_{n}{\bf v}_{n},$ (40)
where ${\bf v}_{n}$ is the $n$-th eigenvector of ${\bf R}_{\rm BS}$ and $\lambda_{n}\geq 0$ is the corresponding eigenvalue. According to (39), we define a compensation matrix ${\bf D}_{\rm BS}$ given by
@@
3:  Generate ${\bf R}_{\rm BS}$ and ${\bf R}_{\rm UE}$ with DPSS according to (42).
4:  Perform EVD on ${\bf R}_{\rm BS}=\mathbf{V}\bm{\Lambda}\mathbf{V}^{-1}$ and ${\bf R}_{\rm UE}=\mathbf{U}{\bm{\Lambda}^{\prime}}\mathbf{U}^{-1}$.
5:  Compensate the phase shift ${\bf V}^{c}={\bf D}_{\rm BS}{\bf V}$, ${\bf U}^{c}\!=\!{\bf D}_{\rm UE}{\bf U}$.
@@
Note that extracting ${\bf D}_{\rm BS}$ from ${\bf R}_{\rm BS}$ only alters the phase of each eigenvector since $(\overline{\mathbf{R}}_{\rm BS}+{\bm{\Gamma}}){\bf D}_{\rm BS}{\bf v}_{n}=\lambda_{n}{\bf D}_{\rm BS}{\bf v}_{n}$. Furthermore, (39) yields
@@
where $L_{\rm UE}=(N_{\rm UE}-1)\lambda/2$ denotes the aperture of the UE array with half-wavelength antenna spacing. Note that $(b)$ holds asymptotically when $N_{\rm UE}$ is sufficiently large. Also, $\overline{\bf R}_{\rm BS}$ is a Toeplitz matrix with each column composed of a shifted sinc function, and the $n$-th eigenvector ${\bf D}_{\rm BS}{\bf v}_{n}$ of this matrix is called the $(n-1)$-th order discrete prolate spheroidal sequence within frequency $W=\kappa L_{\rm UE}/(4\pi\hat{y}^{(0)})$ [37].
Remark 2 ().
@@ V-B Proposed Multipath Near-Field Channel Estimation
denotes the remaining columns of ${\bf G}_{\rm BS}\otimes{\bf G}_{\rm UE}$ under the column selection of ${\rm vec}(\overline{\bm{\Lambda}})$, and can be readily obtained by the estimated coordinates $\{\hat{\bf r}_{\rm S}^{(\ell)}\}_{\ell=1}^{L}$. Substituting (4) and (44) into (9), the measurement model can then be reformulated as
@@ VI Simulation Results
Baseband Samples Computational Complexity (Multiplications) Average Runtime (s)
@@
Greedy $N_{\rm BS}/2$ $N_{\rm BS}G_{x}G_{y}+N_{\rm BS}L(L+1)(2L+1)/6+L^{2}(L+1)^{2}/4$ $0.0548$ $0.1984$ $0.4232$
Proposed 00 $G_{y}\lceil\log_{2}G_{x}\rceil 2^{\lceil\log_{2}G_{x}\rceil}$ $\bf 0.0022$ $\bf 0.0042$ $\bf 0.0065$
@@ VI-A Simulation Setup
Throughout the simulation, we consider the arrays at the BS and UE are equipped with $N_{\rm BS}=256$ and $N_{\rm UE}=4$ antennas with half-wavelength spacing, respectively, and the carrier frequency is set as $f_{c}=28\,{\rm GHz}$. The BS array is placed symmetrically on the $x$-axis and the UE is in the near-field region of the BS as shown in Fig. 1. Unless otherwise specified, we deploy a single RF chain at both the BS and UE. The distances from the UE and $L=5$ scatterers to the center of the BS array are selected uniformly from the a rectangular area limited by $x_{L}=-5\,{\rm m}$, $x_{H}=5\,{\rm m}$, $y_{L}=2\,{\rm m}$, and $y_{H}=25\,{\rm m}$ within the near-field region.
@@ VI-B Power Sensor-based Localization
In this subsection, we compare the performance of the proposed time inversion algorithm with the widely-adopted MUSIC algorithm [22] and greedy grid search [12]. For both baseline algorithms, the number of baseband samples is $N_{\rm BS}/2$, and the near-field region is divided into the same numbers of grids. Particularly, the number of searching grids on the $x$-axis is set as $G_{x}=\eta N_{\rm BS}$, while the searching grids on the $y$-axis are set as $G_{y}=20\eta$, where $\eta\in\{1,2,3\}$ denotes the oversampling rate of searching grids. In contrast, owing to the FFT implementation of the proposed algorithm, the number of searching grids on the $x$-axis are fixed as $G_{x}=\lceil 2(x_{H}-x_{L})/(N_{\rm BS}\lambda)\rceil N_{\rm BS}\approx 8N_{\rm BS}$, while is identical to the baselines on the $y$-axis as $G_{y}=20\eta$. As can be observed from Fig. 5(a), the cumulative distribution functions (CDFs) of localization errors $\epsilon=\sum_{\ell=0}^{L}\|\hat{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm S}^{(\ell)}\|/L$ achieved by three algorithms participating in the comparison are comparable. However, the proposed time inversion algorithm entails two practical advantages over baseline schemes.
@@
First, note that the two baseline algorithms both require $128$ baseband samples, which incurs huge sampling overhead and power consumption in XL-MIMO systems. In contrast, the proposed time-inversion localization algorithm does not need any baseband sampling via high-frequency ADCs. Instead, as mentioned in Section II-B, acquiring power sampling via power sensors is more energy-efficient and economical. Second, the computational complexity of the proposed time inversion algorithm is far lower than that of the two baselines, as is listed in Table I. For instance, according to the values of $G_{x}$ and $G_{y}$ set for different localization algorithms in this subsection, their specific computational complexity can be approximated by $\mathcal{O}(\eta^{2}N_{\rm BS}^{4})$, $\mathcal{O}(\eta^{2}N_{\rm BS}^{2})$, and $\mathcal{O}(\eta N_{\rm BS}\log(N_{\rm BS}))$, respectively. It clearly shows that the proposed algorithm has the lowest complexity among the comparison algorithms even though it employs the largest number of grids on the $x$-axis, and this can also be validated by numerical results of the average runtime in Table I. In particular, the average runtime of the proposed time inversion algorithm is tens and hundreds of times lower than that of the greedy-based grid search and MUSIC algorithm, respectively. Therefore, these two merits of the time inversion algorithm position it an excellent candidate for energy-efficient low-complexity localization algorithm for near-field XL-MIMO systems.
@@ VI-C Sensing-Enhanced CE
Figure 6: (a) The sparse representations of near-field channels under different dictionaries, (b) the mutual correlation matrix $\bm{\Psi}^{H}\bm{\Psi}$ of the spherical wave dictionary and the proposed dictionary, and (c) the approximation error in (42).
@@
  • DFT dictionary: The DFT dictionary is designed according to [8], where the numbers of angle grids are set as $\beta N_{\rm BS}$ and $\beta N_{\rm UE}$ at the BS and UE side, respectively, with $\beta$ being the oversampling rate of the dictionary. Hence, the size of DFT dictionary is $\beta^{2}N_{\rm BS}N_{\rm UE}$.
@@
  • Spherical wave dictionary: Both the distance and angle grids are set as $\lfloor\beta\sqrt{N_{\rm BS}}\rceil$ and $\lfloor\beta\sqrt{N_{\rm UE}}\rceil$ at the BS and UE, respectively. The angle grids are sampled uniformly from $-\pi/2$ to $\pi/2$, while the reciprocal of distance grids are uniformly sampled from $[1/y_{H},1/y_{L}]$ [13]. Therefore, the size of the spherical wave dictionary is $\lfloor\beta\sqrt{N_{\rm UE}}\rceil\lfloor\beta\sqrt{N_{\rm BS}}\rceil\approx\beta^{2}N_{\rm BS}N_{\rm UE}$.
@@
In contrast, the size of the proposed DPSS-based dictionary derived in Section V is $N_{\rm BS}N_{\rm UE}+L$. Additionally, the compressive ratio (CR) of CS-based CE problem is defined as $\mu=\tau/(N_{\rm UE}N_{\rm BS})$ and therefore the required number of baseband samples is $\lfloor\tau N_{\rm BS}^{\rm RF}\rceil=\lfloor\mu N_{\rm BS}^{\rm RF}N_{\rm UE}N_{\rm BS}\rceil$. For fair comparison, the performance of CE achieved by all dictionaries is evaluated based on the OMP algorithm.
@@
We firstly investigate the sparsification performance of the proposed dictionary. The channel sparse representations of the DFT dictionary, spherical wave dictionary, and proposed dictionary are compared in Fig. 6(a), where the sparse representation is obtained by $\tilde{\bf h}=\bm{\Psi}^{\dagger}{\bf h}$. As can be observed, the conventional DFT dictionary shows a severe energy leakage effect in the near-field region [13], which can be improved by the spherical wave dictionary sampled in the polar-domain. Meanwhile, the proposed DPSS-based eigen-dictionary entails the sparsest pattern among the three dictionaries. Note that the proposed dictionary is compensated by the matrix ${\bf D}_{{\rm BS}({\rm UE})}$ in (41) and therefore, the sparse representation contains no specific angular information. In addition, different from the DFT and spherical dictionaries, the support appears in the first several indices since the SVD always sorts the non-zero singular values first.
On the other hand, Fig. 6(b) plots the grayscale map that represents the values of $\bm{\Psi}^{H}\bm{\Psi}$ for spherical wave dictionary (left-hand side) and the proposed dictionary (right-hand side). As can be observed, the codewords in the proposed DPSS-based dictionary are strictly orthogonal to each other, which is far beyond the capabilities of the spherical dictionary. This orthogonality not only inherently reduces the dictionary size but also improves the CE performance, which shall be observed later in this section. We further validate the approximation error of the derivation procedure in (42). As is depicted in Fig. 6(c), the auto-correlation curve stands for the absolute value of $\overline{\bf R}_{\rm BS}[1,:]$, while the red circles show the value of the normalized sinc function. The approximation procedure shows negligible error, which confirms the high accuracy of our proposed approximation $(b)$ in (42).
@@
We then investigate the CE accuracy performance with CR $\mu\in\{0.25,0.4,0.6\}$. The oversampling rate $\beta$ is set to $1$ to keep the sizes of the three considered dictionaries identical. As shown in Fig. 7, both the schemes with the proposed and the spherical wave dictionaries start from lower estimation errors than the DFT dictionary, which validates the alignment of the proposed eigen-dictionary and the near-field spherical wave model. Starting from around $I=10$ iterations, the convergence of the spherical wave dictionary experiences a deceleration due to the increased mutual correlation among codewords, while the other dictionaries converge continuously thanks to the column-wise orthogonality. Additionally, thanks to the accurate localization information incorporated in the DPSS-based dictionary, the proposed CE scheme shows the highest estimation accuracy. Recall that the number of baseband samples is equal to $\lfloor\mu N_{\rm BS}^{\rm RF}N_{\rm UE}N_{\rm BS}\rceil$, which is proportional to $\mu$. Hence, this observation implies that our proposal is more efficient in CE when utilizing the same number of ADC resources, e.g., $410$ samples when $\mu=0.4$, at the baseband. Finally, the achievable NMSE of the proposed method is almost identical with varying CRs $\mu$. This phenomenon indicates that the proposed sensing-enhanced CE has great potential to alleviate the burden of baseband sampling in CE. In other words, in practical systems, only a small number of baseband samples, i.e., $256$, is required to achieve a satisfactory performance in terms of NMSE.

### v3

@@ IV-B Low-Complexity Time Inversion Algorithm
where $\mathcal{F}_{a}[\cdot]$ denotes an FFT operation with padding length $2^{\lceil\log_{2}a\rceil}$, and $\mathcal{F}^{-1}_{a}[\cdot]$ denotes an inverse FFT (IFFT) operation with trimming length $a$. By resorting to FFT, the computational complexity can be greatly reduced from $\mathcal{O}(N_{\rm BS}G_{x})$ to $\mathcal{O}(P\log P)$, where $P=2^{\lceil\log_{2}G_{x}^{\prime}\rceil}$ is the nearest power-$2$ number larger than $G_{x}^{\prime}$.
0:  The range of near-field region $x_{L}$, $x_{H}$, $y_{L}$, and $y_{H}$, the numbers of grids $G_{x}$ and $G_{y}$, multipaths $L$,777 To obtain the value of $L$, an empirical threshold can be adopted to identify the peaks in $\bm{\Theta}$ that are significantly higher than other ripples.  and antennas $N_{\rm BS}$. Measured power pattern $\bf p$ and reference wave ${\bf e}_{\rm ref}$.
@@ V-A Eigen-Dictionary Design for Sparse Representation
${\bf R}_{\rm BS}$ $\overset{~{}~{}~{}}{=}\mathbb{E}\left[{\bf H}{\bf H}^{H}\right]$ (32)
$\overset{~{}~{}~{}}{=}{\bf H}_{\rm LoS}{\bf H}_{\rm LoS}^{H}+\mathbb{E}\left[{\bf H}_{\rm NLoS}{\bf H}_{\rm NLoS}^{H}\right]$
$\overset{~{}~{}~{}}{=}{\bf H}_{\rm LoS}{\bf H}_{\rm LoS}^{H}+\bm{\Gamma},$
@@
${\bf R}_{\rm BS}{\bf u}_{n}={\bf D}_{\rm BS}^{-1}\left(\overline{\mathbf{R}}_{\rm BS}+{\bm{\Gamma}}\right){\bf D}_{\rm BS}{\bf u}_{n}=\lambda_{n}{\bf u}_{n},$ (36)
where ${\bf u}_{n}$ is the $n$-th eigenvector of ${\bf R}_{\rm BS}$ and $\lambda_{n}\geq 0$ is the corresponding eigenvalue. According to (35), we define a compensation matrix ${\bf D}_{\rm BS}$ given by
@@
3:  Generate $\overline{\bf R}_{\rm BS}$ and $\overline{\bf R}_{\rm UE}$ with DPSS according to (38).
4:  Perform EVD on $\overline{\bf R}_{\rm BS}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{-1}$ and $\overline{\bf R}_{\rm UE}=\mathbf{V}{\bm{\Lambda}^{\prime}}\mathbf{V}^{-1}$.
5:  Compensate the phase shift ${\bf U}^{c}={\bf D}_{\rm BS}^{-1}{\bf U}$, ${\bf V}^{c}\!=\!{\bf D}_{\rm UE}^{-1}{\bf V}$.
@@
Note that extracting ${\bf D}_{\rm BS}$ from ${\bf R}_{\rm BS}$ only alters the phase of each eigenvector since $(\overline{\mathbf{R}}_{\rm BS}+{\bm{\Gamma}}){\bf D}_{\rm BS}{\bf u}_{n}=\lambda_{n}{\bf D}_{\rm BS}{\bf u}_{n}$. Furthermore, (35) yields
@@
where $L_{\rm UE}=(N_{\rm UE}-1)\lambda/2$ denotes the aperture of the UE array with half-wavelength antenna spacing. Note that $(b)$ holds asymptotically when $N_{\rm UE}$ is sufficiently large. Also, $\overline{\bf R}_{\rm BS}$ is a Toeplitz matrix with each column composed of a shifted sinc function, and the $n$-th eigenvector ${\bf D}_{\rm BS}{\bf u}_{n}$ of this matrix is called the $(n-1)$-th order discrete prolate spheroidal sequence within frequency $W=\kappa L_{\rm UE}/(4\pi\hat{y}^{(0)})$ [42].
Remark 3.
@@ V-B Proposed Multipath Near-Field Channel Estimation
denotes the remaining columns of ${\bf G}_{\rm BS}\otimes{\bf G}_{\rm UE}$ under the column selection of ${\rm vec}(\overline{\bm{\Lambda}})$, and can be readily obtained by the estimated coordinates $\{\hat{\bf r}_{\rm S}^{(\ell)}\}_{\ell=1}^{L}$.
Substituting (4) and (40) into (9), the measurement model can then be reformulated as
@@ VI Simulation Results
Baseband Samples Computational Complexity (Number of Multiplications) Average Runtime (s)
@@
BF $N_{\rm BS}/2$ $N_{\rm BS}G_{x}G_{y}+N_{\rm BS}L(L+1)(2L+1)/6+L^{2}(L+1)^{2}/4$ $0.0548$ $0.1984$ $0.4232$
STT* $N_{\rm BS}N_{\rm UE}T_{s}$ $8T_{a}\left(N_{\rm BS}+N_{\rm UE}\right)+2N_{\rm BS}G_{x}G_{y}$- -- --
Proposed 00 $G_{y}\lceil\log_{2}G_{x}^{\prime}\rceil 2^{\lceil\log_{2}G_{x}^{\prime}\rceil}$ $\bf 0.0022$ $\bf 0.0042$ $\bf 0.0065$
    $T_{a}$ denotes the number of training iterations.
@@ VI-A Simulation Setup
Throughout the simulation, we consider the arrays at the BS and UE are equipped with $N_{\rm BS}=256$ and $N_{\rm UE}=4$ antennas with half-wavelength spacing [^8] Antenna spacing no greater than half a wavelength effectively avoids $2\pi$ phase wrapping effect., respectively, and the carrier frequency is set as $f_{c}=28\,{\rm GHz}$. The BS array is placed symmetrically on the $x$-axis and the UE is in the near-field region of the BS as shown in Fig. 1. Unless otherwise specified, we deploy a single RF chain at both the BS and UE. The distances from the UE and $L=5$ scatterers to the center of the BS array are selected uniformly from a rectangular area limited by $x_{L}=-5\,{\rm m}$, $x_{H}=5\,{\rm m}$, $y_{L}=2\,{\rm m}$, and $y_{H}=25\,{\rm m}$ within the near-field region. The thermal noise power is set as $\sigma_{\rm n}^{2}=-101\,{\rm dBm}$.
@@ VI-B Power Sensor-based Localization
In this subsection, we compare the performance of the proposed time inversion algorithm with the brute force (BF) grid search, MUSIC algorithm [23], and STT method [24]. The numbers of baseband samples for the three algorithms are set as $N_{\rm BS}/2$, $N_{\rm BS}/2$, and $N_{\rm BS}N_{\rm UE}T_{s}$, respectively, where $T_{s}$ denotes the number of ping-pong pilot iterations for the STT method, and the near-field region is divided into the same number of grids. Particularly, the number of searching grids on the $x$-axis is set as $G_{x}=\eta N_{\rm BS}$, while the searching grids on the $y$-axis are set as $G_{y}=20\eta$, where $\eta\in\{1,2,3\}$ denotes the oversampling rate of searching grids. In contrast, owing to the FFT implementation of the proposed algorithm, the number of searching grids on the $x$-axis is fixed as $G_{x}=\lceil 2(x_{H}-x_{L})/(N_{\rm BS}\lambda)\rceil N_{\rm BS}\approx 8N_{\rm BS}$, while is identical to the baselines on the $y$-axis as $G_{y}=20\eta$. As can be observed from Fig. 5(a), the cumulative distribution functions (CDFs) of localization errors $\epsilon=\sum_{\ell=0}^{L}\|\hat{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm S}^{(\ell)}\|/L$ achieved by four algorithms participating in the comparison are comparable. However, the proposed time inversion algorithm entails two practical advantages over baseline schemes.
@@
First, note that all three baseline algorithms require baseband samples, which incurs huge sampling overhead and power consumption in XL-MIMO systems. In contrast, the proposed time-inversion localization algorithm does not need any baseband sampling via high-frequency ADCs. Instead, as mentioned in Section II-B, acquiring power sampling via power sensors is more energy-efficient and economical. Second, the computational complexity of the proposed time inversion algorithm is far lower than that of the two baselines, as is listed in Table I. For instance, according to the values of $G_{x}$ and $G_{y}$ set for different localization algorithms in this subsection, their specific computational complexity is given by $\mathcal{O}(\eta^{2}N_{\rm BS}^{4})$, $\mathcal{O}(\eta^{2}N_{\rm BS}^{2})$, $\mathcal{O}(\eta^{2}N_{\rm BS}^{2})$, and $\mathcal{O}(\eta N_{\rm BS}\log(N_{\rm BS}))$, respectively. It clearly shows that the proposed algorithm has the lowest complexity among the comparison algorithms even though it employs the largest number of grids on the $x$-axis, and this can also be validated by numerical results of the average runtime in Table I. In particular, the average runtime of the proposed time inversion algorithm is tens and hundreds of times lower than that of the BF grid search and MUSIC algorithm, respectively. Therefore, these two merits of the time inversion algorithm position it an excellent candidate for energy-efficient low-complexity localization algorithm for near-field XL-MIMO systems.
@@ VI-C Sensing-Enhanced CE
Figure 6: (a) The sparse representations of near-field channels under different dictionaries and (b) the approximation error in (38).
@@
  • DFT dictionary: The DFT dictionary is designed according to [8], where the numbers of angle grids are set as $\beta N_{\rm BS}$ and $\beta N_{\rm UE}$ at the BS and UE side, respectively, with $\beta$ being the oversampling rate of the dictionary. Hence, the size of the DFT dictionary is $\beta^{2}N_{\rm BS}N_{\rm UE}$.
@@
  • Spherical wave dictionary: The numbers of grids for distance and angle are set as $\lfloor\beta\sqrt{N_{\rm BS}}\rceil$ and $\lfloor\beta\sqrt{N_{\rm UE}}\rceil$ at the BS and UE, respectively. The angle grids are selected evenly from $-\pi/2$ to $\pi/2$, while the reciprocal of distance grids at angle $\theta$ are evenly selected from $[(1-\cos^{2}\theta)/\sqrt{x_{H}^{2}+y_{H}^{2}},(1-\cos^{2}\theta)/y_{L}]$ [13]. Therefore, the size of the spherical wave dictionary is $(\lfloor\beta\sqrt{N_{\rm BS}}\rceil\lfloor\beta\sqrt{N_{\rm UE}}\rceil)^{2}\approx\beta^{4}N_{\rm BS}N_{\rm UE}$.
@@
In contrast, the size of the proposed DPSS-based dictionary derived in Section V is $N_{\rm BS}N_{\rm UE}+L$. Additionally, the compression ratio (CR) of CS-based CE problem is defined as $\mu=\tau/(N_{\rm UE}N_{\rm BS})$ and therefore the required number of baseband samples is $\lfloor\tau N_{\rm BS}^{\rm RF}\rceil=\lfloor\mu N_{\rm BS}^{\rm RF}N_{\rm UE}N_{\rm BS}\rceil$. For fair comparison, the performance of CE achieved by all dictionaries is evaluated based on the OMP algorithm.
@@
We firstly investigate the sparsification performance of the proposed dictionary. The channel sparse representations of the DFT dictionary, spherical wave dictionary, and proposed dictionary are compared in Fig. 6(a), where the sparse representation is obtained by $\tilde{\bf h}=\bm{\Psi}^{\dagger}{\bf h}$. As can be observed, the conventional DFT dictionary shows a severe energy leakage effect in the near-field region [13], which can be improved by the spherical wave dictionary sampled in the polar-domain. Meanwhile, the proposed DPSS-based eigen-dictionary entails the sparsest pattern among the three dictionaries. Note that the proposed dictionary is compensated by the matrix ${\bf D}_{{\rm BS}({\rm UE})}$ in (37) and therefore, the sparse representation contains no specific angular information. In addition, different from the DFT and spherical dictionaries, the non-zero support set appears in the first several indices since the SVD always sorts the non-zero singular values first. We further validate the approximation error of the derivation procedure in (38). As is depicted in Fig. 6(c), the auto-correlation curve stands for the absolute value of $\overline{\bf R}_{\rm BS}[1,:]$, while the red circles show the value of the normalized sinc function. The approximation procedure shows negligible error, which confirms the high accuracy of our proposed approximation $(b)$ in (38).
Figure 7: Performance comparison of near-field CE error versus pilot length $\tau$ and dictionary oversampling rate $\beta$ when $I=30$.
@@
We then investigate the CE accuracy performance with pilot overhead $\tau$. As shown in Fig. 7, the NMSE performance of CE converges as the pilot length $\tau$ increases and the proposed schemes achieve a lower estimation error than baseline algorithms. Besides, the proposed method converges at around $\tau=300$, while the other baselines converge at around $\tau=500$. Furthermore, to achieve a CE NMSE of $-10\,{\rm dB}$, the proposed method saves up to $85.7\%$ of the pilot overhead when compared to the spherical dictionary with $\beta=1$. In addition, to achieve a CE error of $-15\,{\rm dB}$, the proposed method saves up to $50\%$ and $85.7\%$ of the pilot overhead compared to DFT and spherical dictionaries, respectively, with $\beta=3$. These results clearly demonstrate that the proposed schemes not only efficiently reduce the pilot overhead but also achieve the best estimation error with the aid of the sensing phase.
We next investigate the CE performance against the number of iterations $I$. With the reference of Fig. 7 and recall that $\mu=\tau/N_{\rm BS}N_{\rm UE}$, we set $\mu\in\{0.1,0.3,0.5\}$ for further comparisons. We firstly set the oversampling rate $\beta$ to $1$ to keep the sizes of the three considered dictionaries identical. As shown in Fig. 8, both the schemes with the proposed and the spherical wave dictionaries start from lower estimation errors than the DFT dictionary, which validates the alignment of the proposed eigen-dictionary and the near-field spherical wave model. For cases with $\mu=0.1$, the under-determined observation causes severe upward warping phenomenon as shown at the right side of the curves, while the proposed method still achieves satisfying NMSE performance. For cases with $\mu>0.1$, the convergence of the spherical wave dictionary experiences a deceleration due to the insufficient distance sampling interval, while the other dictionaries converge continuously. Additionally, thanks to the accurate localization information incorporated in the DPSS-based dictionary, the proposed CE scheme shows the highest estimation accuracy. Recall that the number of baseband samples is equal to $\lfloor\mu N_{\rm BS}^{\rm RF}N_{\rm UE}N_{\rm BS}\rceil$, which is proportional to $\mu$. Hence, this observation implies that our proposal is more efficient in CE when utilizing the same number of ADC resources, e.g., $307$ samples when $\mu=0.3$, at the baseband. Finally, the achievable NMSE of the proposed method is almost identical with varying CRs $\mu$. This phenomenon indicates that the proposed sensing-enhanced CE has great potential to alleviate the burden of baseband sampling in CE. In other words, in practical systems, only a small number of baseband samples, i.e., $0.1N_{\rm BS}N_{\rm UE}$, is required to achieve a satisfactory performance in terms of NMSE.